Problem 2 If the cost of electricity in your area is 12 cents per kWh, what is the total cost (capital plus operation) of using a 60-W incandescent bulb for 1000 hours (the lifetime of that bulb) if the bulb cost 25 cents? If we replace this bulb with a compact fluorescent light that provides the same light output, but at one- quarter the wattage, and which costs $1.50 but lasts 10 times longer (10,000 hours), what will that total cost be? How do they compare?
Answer #1

For bulb:

energy used = Power x time

= 60 x 1000 = 60,000 Wh

= 60 kWh

energy cost = 60 x 12 = 720 cents  

total cost = 720 + 25 = 745 cents Or 7.45 dollars

For CFL,

energy used = 25 x 10,0000 = 250 kWh

energy cost = 250 x 12 = 3000 cents Or 30 dollars

total cost = 30 + 1.50 = 31.5 dollars

Comparision: for 1000 hrs, cost of CFL is 3.15 dollars.

that is less than half of cost of bulb.

hence using CFL is economic.